OpenStudy (anonymous):

In how many ways can the letters A, T, S, I, N, and O be ordered? A. 24 B. 120 C. 720 D. 5,040

OpenStudy (anonymous):

B. 120

OpenStudy (anonymous):

\[5 \times 4 \times 3 \times 2 \times 1\]

OpenStudy (anonymous):

@jazleday1

OpenStudy (anonymous):

Thanks :)